OpenStudy (anonymous):

Solve for x. 7x = 78

OpenStudy (anonymous):

x = 78 / 7

OpenStudy (anonymous):

You want to isolate x on one side of the equation. Since we have 7 times x on the left side, we want to divide both sides by 7. On the left side this will leave you with x by itself.

OpenStudy (uri):

7x = 78 x=78/7 x=11.1428

OpenStudy (anonymous):

so x=11 1/7?

OpenStudy (anonymous):

no just x = 78 / 7

OpenStudy (texaschic101):

or 11 1/7

OpenStudy (anonymous):

how??

OpenStudy (texaschic101):

78/7 is an improper fraction....turning it into a mixed number makes it 11 1/7

OpenStudy (anonymous):

ok what ever you can say what ever you want

OpenStudy (texaschic101):

either answer is correct

hero (hero):

\[\frac{78}{7} = \frac{77 + 1}{7} = \frac{77}{7} + \frac{1}{7} = 11 + \frac{1}{7} = 11\frac{1}{7}\]

hero (hero):

@friedrice

OpenStudy (anonymous):

hi!!

OpenStudy (anonymous):

i already know hero thanks

hero (hero):

I was just showing you why \(\large\frac{78}{7} = 11\frac{1}{7}\)